1. What is the vapor pressure of the pure solvent if the vapor pressure of a solution of 15g sucrose (C12H22O11) in 100g of ethanol (C2H5OH) is 58 mmHg?
1
Expert's answer
2022-02-15T03:02:02-0500

Given vapor pressure of the solution,Psolution = 55 mmHg

Given the mass of sucrose (C6H12O6) = 10 g

molar mass of sucrose (C6H12O6) = 180 g/mol

Hence moles of sucrose in the solution = mass / molar mass = 10 g / 180 g = 0.05556 mol

mass of ethanol = 100 g

molar mass of ethanol = 46 g/mol

Hence moles of ethanol = mass / molar mass = 100 g / 46 g/mol = 2.174 mol

Hence mole fraction of solvent, ethanol, Xethanol = 2.174 mol / (2.174 + 0.05556) mol = 0.975

Now applying Raoult's Law

Psolution = Xethanol*P0ethanol

=> P0ethanol = Psolution / Xethanol  = 55 mmHg / 0.975 = 56.4 mm Hg